# Nonlinear Superconducting Magnetoelectric Effect

### v4

@@ Abstract
To illustrate NSM effect in the ALM/SC heterostructure, we numerically study the supercurrent induced magnetization in Fig. 2 for the Bloch Hamiltonian in Eq. (8). By applying the current along $x$ direction as $\bm{q}=(q_{x},0)$, we calculate the magnetization directly from Eq. (4) at $J_{\mathrm{ex}}=0.2t$ as shown in Fig. 2(b). One finds that $M_{z}\sim q_{x}^{2}$ as expected from Fig. 1(b). It is also worth noting that when $q_{x}=0$, equilibrium magnetization vanishes $M_{z}=0$, which is consistent with $C_{4}\mathcal{T}$ symmetry in altermagnets. We note that numerically computing the NSM susceptibility $\chi_{xx}^{z}$ from Eq. (7) and multiplying by $q_{x}^{2}$ as plotted in the dashed green line matches the solid line in orange [from Eq. (4)].
In Fig. 2(c), we show the second-order spin susceptibility $\chi_{xx}^{z}$ as a function of $J_{\mathrm{ex}}$ from Eq. (7). Generally, $\chi_{xx}^{z}$ starts to grow linearly with $J_{\mathrm{ex}}$ as shown in the zoom-in inset of Fig. 2(c). Furthermore, we also calculate the first-order spin susceptibility $\alpha_{xy}$ (purple) which is shown in Fig. 2(d). As expected, $\alpha_{xy}$ directly depends on $\lambda$; for small $\lambda$, $\alpha_{xy}$ increases linearly with $\lambda$ and vanishes when $\lambda=0$. In contrast, for NSM $\chi_{xx}^{z}$ is finite for $\lambda=0$ and exhibits a weak dependence on $\lambda$; $\chi_{xx}^{z}$ becomes smaller when $\lambda$ is nonzero since Rashba SOI tends to pin the spin in the in-plane direction and weakens the out-of-plane spin polarization. In our calculations, we adopt the BCS temperature dependence (for the $s$-wave superconductor substrate) of $\Delta_{0}$ with $\Delta(T)=\Delta_{0}\tanh(1.74\sqrt{T_{c}/T-1})$ and $\Delta_{0}=1.76k_{B}T_{c}$.
Figure 3: The second-order spin susceptibility $\chi_{xx}^{z}$ as a function of the temperature $T$ at $(J_{\mathrm{ex}},\Delta_{0},\mu)=(0.02,0.1,0.3)t$. The purple diamonds denotes the numerical results from Eq. (7) and the orange line denotes the analytical result from Eq. (9). At low temperatures, $\chi^{z}_{xx}\sim\beta^{3}\mathrm{sech}^{2}\beta$.
@@
$\begin{split}f(x)=&\frac{\beta^{3}J_{\mathrm{ex}}}{64\pi}x\mathrm{sech}^{2}\gamma\Big{[}\frac{4}{\beta^{2}}+t^{2}x^{4}(2-3\mathrm{sech}^{2}\gamma)\\ &+\frac{4tx^{2}(\mu-tx^{2})\tanh\gamma}{\gamma}\Big{]},\end{split}$ (9)
where $\gamma=\beta\sqrt{\Delta(T)^{2}+(tx^{2}-\mu)^{2}}/2$. By employing Eq. (9), we plot the $\chi_{xx}^{z}$ as a function of temperature in Fig. 3 (orange solid curve) with $J_{\mathrm{ex}}=0.02t$. By way of comparison, we also show a fully numerical plot of Eq. (7) (purple diamonds); both agree with each other. Interestingly, $\chi_{xx}^{z}$ vanishes at zero temperature. At low temperatures ($\beta\Delta_{0}\gg 1$), $\chi_{xx}^{z}$ starts to grows with $T$ scaling as $\chi_{xx}\sim\beta^{3}\mathrm{sech}^{2}\beta$. However, at $T/T_{c}\sim 0.6$, it reaches a peak after which it rapidly diminishes with temperature. This non-monotonic behavior contrasts sharply with that of the linear magnetoelectric effect in Rashba-type superconductors He et al. (2019); Not that instead saturates at low temperature.
In superconductors, supercurrent flow can arise due to either an external applied magnetic field or an explicit transport current Anthore et al. (2003). In the former case, there is a screening supercurrent on the surface of the superconductor. In the latter case, the global excitation of the superconductor gives a uniform Cooper pair momentum $\bm{q}$. Specifically, for slowly varying AC supercurrent as $\bm{J}_{s}(t)=\bm{J}_{\omega}\cos\omega t$ with $\omega\ll\Delta_{0}$, we expect an oscillatory Cooper pair momentum $\bm{q}(t)=\bm{q}_{\omega}\cos\omega t$. This will allow to obtain a second harmonic magnetization in the weak oscillating regime with $\chi_{ab}^{c}(2\omega)=\chi_{ab}^{c}/4$.
@@
Candidate materials.—We anticipate that NSM we discuss here can be realized in ALM/SC formed out of readily available $d$-wave altermagnetic materials Šmejkal et al. (2022a). As an example, consider the altermagnet RuO2 with in-plane lattice constant $a=4.5\AA$, and $t=2.5$ eV, $J_{\mathrm{ex}}=0.5$ eV Šmejkal et al. (2022b); Zhang et al. (2024). Using a pairing gap of $\Delta_{0}=1$ meV, $\mu=10$ meV, $\lambda=10$ meV and $T=0.3T_{c}$ we estimate $\chi_{xx}^{z}\approx 0.07\mu_{B}$. For a Cooper pair momentum $q\approx 10^{-3}\AA^{-1}$ Zhu et al. (2021), the induced out-of-plane spin density in the RuO2 thin film/superconductor heterostructure is $\sim 0.7\times 10^{-5}\mu_{B}/\mathrm{nm}^{2}$. Note that this value is comparable to the spin generation in MnBi2Te4 and other noncentrosymmetric ferromagnetic systems Xiao et al. (2022); Chernyshov et al. (2009); Fang et al. (2011). Another example is KRu4O8 with the parameter values: $t=0.05$ eV, $J_{\mathrm{ex}}=0.018$ eV, and $a=9.9\AA$. We find $\chi_{xx}^{z}\approx 0.01\mu_{B}$ and out-of-plane spin density similar to that discussed above. These magnitudes are estimated qualitatively using a low-energy effective description of altermagnets in the small $J$ and weak doping (small $\mu$) regime as well as a gapped BdG spectrum. In real altermagnets with large $J$ and high doping, the BdG spectrum becomes gapless. We anticipate this may enhance NSM with additional contribution from the quasiparticles excited at finite temperature.
In this work, we have proposed a novel nonlinear superconducting magnetoelectric effect. Naturally occurring in ALM/SC heterostructures, this second-order spin generation in response to the supercurrent becomes the leading order contribution to the magnetization along the out-of-plane direction. Supercurrent induced magnetization can be readily detected by using a superconducting quantum interference device (SQUID) perpendicular to the altermagnet surface, which can be employed to probe the magnetic flux change Chirolli et al. (2022). Interestingly, nonlinear spin generation that is second order in an applied electric field can be realized in normal-state (non-superconducting) altermagnets with $C_{4}\mathcal{T}$ symmetry Xiao et al. (2022). Our work shows that nonlinear spin generation survives in the superconducting state and is driven by a dissipationless supercurrent.
@@
Acknowledgements.—We thank Bo Yang and Wen-Yu He for helpful discussions. This work was supported by the Ministry of Education Singapore under its Academic Research Fund Tier 2 Grant No. MOE-T2EP50222-0011.
Note added.—Recently, we became aware of an independent pre-print posted on the arXiv in Ref. Zyuzin (2024) which also found a second-order supercurrent-induced spin polarization in superconductors with $d$-wave magnetization.
@@
I I. Theory of nonlinear superconducting Magnetoelectric effect
I.1 A. Green’s function method
In this section we provide more details for deriving the general formula of superconducting magnetoelectric effect, including both the linear and nonlinear terms. The superconducting state can be intrinsic or proximity induced with a pairing function $\hat{\Delta}_{\bm{k}}$. $\hat{\Delta}_{\bm{k}}$ can be conventional $s$-wave with a uniform pairing gap $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$ or unconventional pairing states. In the current-carrying state, the finite-$\bm{q}$ Bogoliubov-de Gennes (BdG) Hamiltonian in the Nambu basis $(\hat{c}_{\bm{k}+\bm{q}/2,\uparrow},\hat{c}_{\bm{k}+\bm{q}/2,\downarrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\uparrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\downarrow})$ reads
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\left(\begin{matrix}{}H_{\bm{k}+\bm{q}/2}&\hat{\Delta}_{\bm{k}}\\ \hat{\Delta}^{\dagger}_{\bm{k}}&-H^{*}_{-\bm{k}+\bm{q}/2}\end{matrix}\right)$ (S1)
@@
In the main text, we focus on the altermagnet/superconductor heterostructure. The total free energy $\mathcal{F}$ of the system contains the superconductor part $\mathcal{F}_{\mathrm{sc}}$ and the altermagnet part $\mathcal{F}_{\mathrm{al}}$ with $\mathcal{F}=\mathcal{F}_{\mathrm{sc}}+\mathcal{F}_{\mathrm{al}}$. The supercurrent $J_{s}=\nu_{s}\bm{q}$ with the superfluid density $\nu_{s}=\partial^{2}\mathcal{F}_{\mathrm{sc}}(\bm{q})/\partial^{2}\bm{q}$. The partition function $\mathcal{Z}(\bm{h})$ is given by
@@
$\begin{split}M_{a}&=-\frac{1}{\beta}\frac{\partial}{\partial{h_{a}}}\mathrm{log}Z(\bm{h})\\ &=-\frac{1}{2\beta}\sum_{\bm{k},\bm{q},n}\frac{\partial}{\partial{h_{a}}}\mathrm{Tr}[\mathrm{log}G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})]\end{split}$ (S4)
@@
$\Sigma(\bm{h})=\left(\begin{array}[]{cc}\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}&0\\ 0&-\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}\end{array}\right),$ (S5)
@@
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\mathcal{H}_{\mathrm{BdG}}^{\bm{k}}+q_{a}\hat{v}_{a}/2+q_{a}q_{b}\hat{w}_{ab}/8+\mathcal{O}(\bm{q}^{3})$ (S7)
The velocity operators are
@@
I.2 B. Analytical result
In the previous subsection, we use the Green’s function method to give the general formulas of the linear and nonlinear magnetization. These formulas can be adopt to evaluate the spin susceptibilities for altermagnet/superconductor heterostructures. In this subsection, we provide the analytical solution of the nonlinear spin susceptibility $\chi^{z}_{ab}$. For the altermagnets, we can write down a general two-band model
$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+J_{ex}a^{2}[2k_{x}k_{y}\cos(2\theta)+(k_{x}^{2}-k_{y}^{2})\sin(2\theta)]s_{z}$ (S13)
$t$ parameterizes the usual kinetic energy, $J_{ex}$ denotes the $d$-wave exchange magnetic order parameter and $\mu$ is the chemical potential. $\theta$ is the angle between the altermagnetic orientation and the crystalline axes. The velocity operators are
@@
$\hat{w}_{xx}=a^{2}\left(\begin{array}[]{cc}2ts_{0}+2J_{ex}\sin(2\theta)s_{z}&0\\ 0&-2ts_{0}-2J_{ex}\sin(2\theta)s_{z}\end{array}\right),\hat{w}_{xy}=a^{2}\left(\begin{array}[]{cc}2J_{ex}\cos(2\theta)s_{z}&0\\ 0&-2J_{ex}\cos(2\theta)s_{z}\end{array}\right),$ (S16)

### v5

@@ Abstract
To illustrate NSM effect in the ALM/SC heterostructure, we numerically study the supercurrent induced magnetization in Fig. 2 for the Bloch Hamiltonian in Eq. (7). By applying the current along $x$ direction as $\bm{q}=(q_{x},0)$, we plot the magnetization directly from Eq. (4) at $J_{\mathrm{ex}}=0.2t$ in Fig. 2(b). We find that $M_{z}\sim q_{x}^{2}$ as expected from Fig. 1(b). It is also worth noting that when $q_{x}=0$, equilibrium magnetization vanishes $M_{z}=0$, which is consistent with $C_{4}\mathcal{T}$ symmetry in altermagnets. We note that numerically computing the NSM susceptibility $\chi_{xx}^{z}$ from Eq. (6) and multiplying by $q_{x}^{2}$ as plotted in the dashed green line matches the solid line in orange [from Eq. (4)].
@@
$\begin{split}f(\zeta)=&\frac{\beta^{3}}{64\pi}\zeta\mathrm{sech}^{2}\gamma\Big{[}\frac{4}{\beta^{2}}+t^{2}\zeta^{4}(2-3\mathrm{sech}^{2}\gamma)\\ &+\frac{4t\zeta^{2}(\mu-t\zeta^{2})\tanh\gamma}{\gamma}\Big{]},\end{split}$ (8)
where $\gamma=\beta\sqrt{\Delta(T)^{2}+(t\zeta^{2}-\mu)^{2}}/2$. Furthermore, we also calculate the first-order spin susceptibility $\alpha_{xy}$ (purple) which is shown in Fig. 2(d). As expected, $\alpha_{xy}$ directly depends on $\lambda$; for small $\lambda$, $\alpha_{xy}$ increases linearly with $\lambda$ and vanishes when $\lambda=0$.
The role of $\lambda$ (and more generally other SOI) for NSM $\chi_{xx}^{z}$ is more nuanced. At zero temperature, $\chi_{xx}^{z}$ vanishes as $\chi_{xx}^{z}\sim\beta^{3}\mathrm{sech}^{2}\beta$ [Eq. (8)] when $\lambda=0$ since $[\eta_{z},\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}]=0$ rendering $\langle\eta_{z}\rangle$ conserved. When $\lambda\neq 0$ is turned on, $\langle\eta_{z}\rangle$ is no longer conserved turning on $\chi_{xx}^{z}$. In contrast, at finite temperature $\chi_{xx}^{z}$ is sizeable even for $\lambda=0$ yielding a nonzero $M_{z}$ that arises from thermal fluctuations at finite temperature in the current-carrying state. Indeed, $\chi_{xx}^{z}$ at finite temperature (e.g., $T=0.3T_{\rm c}$) in Fig. 2(d) exhibits a weak dependence on $\lambda$. This evidences the large altermagentic spin splitting Šmejkal et al. (2022a) of order several hundred millielectronvolts captured by $J_{\rm ex}\gg\lambda$ dominates the magnetic order yielding a robust NSM.
NSM in a noncollinear antiferromagnet.—So far we have focused on NSM in $d$-wave altermagnets—a type of collinear antiferromagnet. The general NSM in Eq. (6) can exist in a larger class of magnet/superconductor heterostructures (see e.g., Table 1 as well as SM Not for a detailed discussion). Indeed, based on Table 1, we find that for centrosymmetric noncollinear antiferromagnet with $D_{3d}$ symmetry, an in-plane NSM with $\chi_{xx}^{y}=\chi_{xy}^{x}=\chi_{yx}^{x}=-\chi_{yy}^{y}\neq 0$ can manifest. To capture the physics of a metal with noncollinear magnetic order on a kagome lattice, we adopt the low energy Hamiltonian Lee et al. (2024a); Not as
$H_{\bm{k}}=(ta^{2}\bm{k}^{2}-\mu)s_{0}+ma^{2}[2k_{x}k_{y}s_{x}+(k_{x}^{2}-k_{y}^{2})s_{y}],$ (9)
where $m$ captures the noncollinear magnetic order. Note that the noncollinear magnetic order induced winding spin texture on the Fermi surface [Fig. 3(a)] sharply contrasts with the collinear magnetic order in $d$-wave altermagnets [Fig. 1(c)]. As we now explain, this difference in magnetic order leads to an in-plane NSM.
To see this, we follow the same procedure as described above and utilize the parameters $(m,\mu,\Delta_{0})=(0.2,0.3,0.1)t$. We obtain an in-plane NSM in Fig. 3(b) that dominates the magnetoelectricity; indeed, as expected from Table. 1 (last column), $M_{y}$ is quadratic in $q_{x}$ – it is the leading order in-plane magnetoelectric term. In-plane NSM is directly controlled by the noncollinear magnetic order as evidenced in Fig. 3(c), where $\chi_{xx}^{y}$ turns on with $m$. As a result, we find that NSM arises in the presence of a variety of magnetic orders: such as $J_{\mathrm{ex}}$ (altermagnet, collinear) or $m$ (kagome, noncollinear) yielding both out-of-plane as well as in-plane NSM responses.
Figure 3: (a) The Fermi surface associated with the winding spin texture of Eq. (9). Blue (yellow) color denotes the upper (lower) band. (b) The spin magnetization $M_{y}$ as a function of $q_{x}$ from Eq. (4) (solid orange) aligns with $\chi^{y}_{xx}q_{x}^{2}$ from Eq. (6) (dashed green). (c) $\chi^{y}_{xx}$ as a function of noncollinear magnetic order $m$. We set $(\Delta_{0},\mu)=(0.1,0.3)t$ and $T=0.3T_{c}$.
@@
Discussion.—We expect that NSM can be realized in several concrete materials. As an example, considering the altermagnet RuO2 thin film with lattice constant $a=4.5\AA$, $c=3.1\AA$, and $t=2.5$ eV, $J_{\mathrm{ex}}=0.5$ eV Šmejkal et al. (2022b); Zhang et al. (2024). Using a pairing gap of $\Delta_{0}=1$ meV, $\mu=0.1$ eV, $\lambda=40$ meV and $T=0.3T_{c}$ we estimate $\chi_{xx}^{z}\approx 0.1\mu_{B}$. For a Cooper pair momentum $q\approx 10^{-3}\AA^{-1}$ Zhu et al. (2021), the induced out-of-plane spin density in the RuO2 thin film/superconductor heterostructure is $\sim 0.3\times 10^{-4}\mu_{B}/\mathrm{nm}^{3}$. Note that this value is comparable to the spin generation in MnBi2Te4 and other noncentrosymmetric ferromagnetic systems Xiao et al. (2022); Chernyshov et al. (2009); Fang et al. (2011). Another example is KRu4O8 with the parameter values: $t=0.05$ eV, $J_{\mathrm{ex}}=0.018$ eV, and $a=9.9\AA$ Zhang et al. (2024). We find $\chi_{xx}^{z}\approx 0.01\mu_{B}$ and out-of-plane spin density similar to that discussed above. The induced NSM magnetization densities (of order $\sim 10^{-4}\mu_{B}/\mathrm{nm}^{3}$ corresponding to equivalent magnetic fields $\sim 1\mu{\rm T}$) can be readily detected by superconducting quantum interference devices (SQUID) Chirolli et al. (2022); Vasyukov et al. (2013) or diamond nitrogen-vacancy center-based magnetometers Balasubramanian et al. (2009).
Naturally, high quality interfaces should be used, since disorder can locally disrupt the superconducting pairing and suppress the supercurrent. Additionally, in our analysis above, we have focussed on ideal ALM/SC heterostructures; in realistic systems, surface reconstruction may ensue. For example, reconstruction can produce to interfacial spin-orbit coupling that breaks local mirror symmetry Linder and Yokoyama (2011), inducing additional linear magnetoelectric responses including in the out-of-plane direction (see SM Not of interface effects). Nevertheless, we expect the large altermagnetic spin-splitting Šmejkal et al. (2022a) readily dominates the response rendering NSM robust.
Even in the presence of an out-of-plane linear magnetoelectric response, we anticipate that NSM can be distinguished by dynamical probes: e.g., oscillatory Cooper pair momentum $\bm{q}(t)=\bm{q}\cos\omega t$ for slowly varying AC supercurrent with $\omega\ll\Delta_{0}$. In this weak oscillating regime, a rectified as well as a second harmonic magnetization appear: $\delta M_{c}^{(2)}(t)=\chi^{c}_{ab}q_{a}q_{b}(1+\cos 2\omega t)/2$. Both rectified as well as second harmonic depart from that of the conventional linear magnetoelectric effect, which produces an oscillating magnetization at frequency $\omega$. This distinction is important practically since it enables to directly identify NSM even in systems where NSM is not the leading order magnetoelectric effect: e.g., conventional linear magnetoelectric effect produces a magnetization that averages to zero over one cycle, while the rectified NSM does not vanish allowing to directly identify its contribution; similarly locking-in to $2\omega$ allows to directly distinguish between linear and nonlinear superconducting magnetoelectric effects.
While NSM is naturally realized in ALM/SC heterostructures (driven by its large spin-splitting), we find that it persists more generally for a wide range of other materials and symmetries (see SM Not ). Strikingly, when NSM is combined with the linear magnetoelectric effect, current induced magnetization can become non-reciprocal producing distinct magnetizations for $+\bm{q}$ and $-\bm{q}$ Not , a new tool for non-dissipative spintronics. Perhaps most interesting, however, is the NSM’s ability to distinguish dilineate magnetic orders: as a result, we expect NSM provides a promising approach for exploring the rich interplay between magnetism and superconductivity in unconventional magnetic materials that include kagome materials Yin et al. (2022), as well as materials in different spin space groups Xiao et al. (2024); Jiang et al. (2024); Chen et al. (2024).
@@
Acknowledgements.—We thank Bo Yang, Wen-Yu He, Naizhou Wang and Benjamin T. Zhou for helpful discussions. This work was supported by the Ministry of Education Singapore under its Academic Research Fund Tier 2 Grant No. MOE-T2EP50222-0011.
@@
S1 Microscopic theory of superconducting magnetoelectric effect
S1.1 Green’s function method
We provide more details in deriving the general formula of superconducting magnetoelectric effect, including both the linear and nonlinear terms. The superconducting state can be intrinsic or proximity induced with a pairing function $\hat{\Delta}_{\bm{k}}$. $\hat{\Delta}_{\bm{k}}$ can be conventional $s$-wave with a uniform pairing gap $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$ or unconventional pairing states. In the current-carrying state, the finite-$\bm{q}$ mean field Hamiltonian is $H_{\mathrm{mf}}(\bm{q})=\frac{1}{2}\Psi_{\bm{k},\bm{q}}^{\dagger}\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}\Psi_{\bm{k},\bm{q}}$ in the Nambu basis $\Psi_{\bm{k},\bm{q}}=(\hat{c}_{\bm{k}+\bm{q}/2,\uparrow},\hat{c}_{\bm{k}+\bm{q}/2,\downarrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\uparrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\downarrow})^{T}$. The Bogoliubov-de Gennes (BdG) Hamiltonian reads
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\left(\begin{matrix}{}H_{\bm{k}+\bm{q}/2}&\hat{\Delta}_{\bm{k}}\\ \hat{\Delta}^{\dagger}_{\bm{k}}&-H^{*}_{-\bm{k}+\bm{q}/2}\end{matrix}\right),$ (S1)
@@
In the main text, we focus on the altermagnet/superconductor heterostructure. Therefore the total free energy $\mathcal{F}$ of the system contains the superconductor part $\mathcal{F}_{\mathrm{sc}}$ and the material part $\mathcal{F}_{m}$ with $\mathcal{F}=\mathcal{F}_{\mathrm{sc}}+\mathcal{F}_{\mathrm{m}}$. The supercurrent $J_{s}=\nu_{s}\bm{q}$ with the superfluid density $\nu_{s}=\partial^{2}\mathcal{F}_{\mathrm{sc}}(\bm{q})/\partial^{2}\bm{q}$. The partition function $\mathcal{Z}(\bm{h})$ is given by
@@
$\begin{split}M_{a}&=-\frac{1}{\beta}\frac{\partial}{\partial{h_{a}}}\mathrm{log}Z(\bm{h})\\ &=-\frac{1}{2\beta}\sum_{\bm{k},\bm{q},n}\frac{\partial}{\partial{h_{a}}}\mathrm{Tr}[\mathrm{log}G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})].\end{split}$ (S4)
@@
$\Sigma(\bm{h})=\left(\begin{array}[]{cc}\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}&0\\ 0&-\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}\end{array}\right).$ (S5)
@@
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\mathcal{H}_{\mathrm{BdG}}^{\bm{k}}+q_{a}\hat{v}_{a}/2+q_{a}q_{b}\hat{w}_{ab}/8+\mathcal{O}(\bm{q}^{3}),$ (S7)
where the velocity operators are
@@
S1.2 Analytical result
In the previous subsection, we use the Green’s function method to give the general formulas of the linear and nonlinear magnetization. These formulas can be adopted to evaluate the spin susceptibilities for altermagnet/superconductor heterostructures. In this subsection, we provide the analytical solution of the nonlinear spin susceptibility $\chi^{z}_{ab}$. For ideal altermagnets, we can write down a general two-band model
$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+J_{ex}a^{2}[2k_{x}k_{y}\cos(2\theta)+(k_{x}^{2}-k_{y}^{2})\sin(2\theta)]s_{z},$ (S13)
where $t$ parameterizes the usual kinetic energy, $J_{ex}$ denotes the $d$-wave exchange magnetic order parameter and $\mu$ is the chemical potential. $\theta$ is the angle between the altermagnetic orientation and the crystalline axes. The velocity operators are
@@
$\hat{w}_{xx}=a^{2}\left(\begin{array}[]{cc}2ts_{0}+2J_{ex}\sin(2\theta)s_{z}&0\\ 0&-2ts_{0}-2J_{ex}\sin(2\theta)s_{z}\end{array}\right),\hat{w}_{xy}=a^{2}\left(\begin{array}[]{cc}2J_{ex}\cos(2\theta)s_{z}&0\\ 0&-2J_{ex}\cos(2\theta)s_{z}\end{array}\right).$ (S16)